OpenStudy (anonymous):

HELP!!!??? integrate e^(6x)*sin(7x) dx

OpenStudy (xapproachesinfinity):

integration by part did you try anything

OpenStudy (anonymous):

I got: \[\int\limits e ^{6x}\sin7x dx + \frac{ 49}{ 36 }\int\limits e ^{6x}\sin7x dx = \frac{ e ^{6x}\sin7x }{ 6 } - \frac{ 7e ^{6x}\cos7x }{ 36 }\]

OpenStudy (xapproachesinfinity):

hmm does not seem right to me

OpenStudy (xapproachesinfinity):

show the work done to get that

OpenStudy (xapproachesinfinity):

part of your solution is good you got problem with number you got in the result

OpenStudy (xapproachesinfinity):

just make sure about what you did again perhaps you can spot a mistake there